### 06 Mar 2026 v1

@@ 5.3 Lorenz–96 model
where $\mathcal{H}$ is a linear subsampling operator that extracts every other state variable,
@@
and the observation noise level is set to $\sigma=1$. This setup represents a partially observed scenario in which only half of the state variables are directly measured.
@@
Figure 11: Reconstructed spatial state fields (Left) and corresponding error (Right) with the observation interval $\Delta t=0.1$ and latent dimension $n=64$. From top to bottom, the results show the localization EnKF, AE-EnKF, DAE-EnKF and the LAE-EnKF.
@@
Figure 12: Time evolution of a representative observed variable (Left) and an unobserved variable (Right) reconstructed with observation interval $\Delta t=0.1$ and latent dimension $n=64$, compared with the corresponding true trajectories. From top to bottom, the results show the localization EnKF, AE-EnKF, DAE-EnKF and the LAE-EnKF.
Dense observation regime
We first consider a relatively dense observation setting with observations available every $\Delta t=0.1$. LAE models are trained with latent dimensions $n\in\{64,80,100\}$, and the resulting LAE-EnKF is compared with AE-EnKF, DAE-EnKF, and an EnKF baseline. Tab. 3 reports the the global relative RMSE $\mathrm{E}_{\mathrm{Rel},1:T}$ over the full state.
For reference, we report results for the standard EnKF both with and without covariance localization. As is well known for the Lorenz–96 system, the EnKF without localization suffers from severe sampling errors and produces extremely poor assimilation results. Without localization, spurious long-range correlations dominate the analysis step, leading to filter instability and large reconstruction errors. For this reason, unless otherwise stated, all subsequent references to EnKF correspond to the localized EnKF implementation. In contrast, all three latent-space methods—AE-EnKF, DAE-EnKF, and the proposed LAE-EnKF—are implemented using the standard EnKF formulation without covariance localization. This choice is intentional and highlights a key distinction: the latent-space representations are designed to mitigate sampling errors through structured dynamics, rather than through ad hoc localization techniques.
Fig. 11 shows the reconstructed spatial fields and the corresponding errors for the representative case $n=64$, while Fig. 12 presents the temporal evolution of one observed and one unobserved state variable. AE-EnKF and DAE-EnKF exhibit larger reconstruction errors and irregular temporal behavior, indicating instability in the learned latent dynamics. In contrast, LAE-EnKF consistently achieves lower errors and smoother temporal evolution for both observed and unobserved variables. This demonstrates that enforcing stable linear dynamics in the latent space substantially improves the recovery of hidden states in a chaotic system.
@@
Figure 13: Reconstructed spatial state fields (Left) and corresponding error (Right) with the observation interval $\Delta t=0.2$ and latent dimension $n=64$. From top to bottom, the results show the localization EnKF, AE-EnKF, DAE-EnKF and the LAE-EnKF.
@@
Sparse observation regime
@@
Table 3: Relative RMSE $\mathrm{E}_{\text{Rel},1:T}$ for different methods with varying latent dimension $n$ under two observation intervals.
$\boldsymbol{n}$ EnKF AE- DAE- LAE-
w/o loc. w/ loc. EnKF EnKF EnKF
$\boldsymbol{\Delta t=0.1}$ - 0.5305 0.1512
$n=64$ 0.4584 0.8837 0.1477
$n=80$ 0.4940 0.7523 0.1512
$n=100$ 0.3998 0.7254 0.1429
$\boldsymbol{\Delta t=0.2}$ 0.3698 0.1697
$n=64$ 0.3733 1.1631 0.1585
$n=80$ 0.3092 0.7142 0.1605
$n=100$ 0.3225 0.6590 0.1531
@@ 6 Conclusion
We denote by $\mathcal{F}_{\mathrm{NN}}(d_{1},d_{2};L,p,K,\kappa,R)$ the class of neural networks introduced in [25], where $d_{1}$ and $d_{2}$ are the input and output dimensions. The networks have depth $L$, width $p$, and at most $K$ nonzero parameters, each bounded by $\kappa$ in magnitude. The effective parameter count $K$ may depend on $(d_{1},d_{2})$. For inputs satisfying $\|\mathbf{x}\|_{\infty}\leq B$, the network outputs are uniformly bounded by $R$. The proof is based on controlling both approximation and estimation errors through a covering number bound for the class $\mathcal{F}_{\mathrm{NN}}$ under the uniform norm. To make the argument precise, we first introduce the assumptions and auxiliary lemmas needed for the subsequent analysis.
Definition A.1 (Covering number [35, Definition 2.1.5]).
@@
Suppose Assumption 1 holds. For any $0<\varepsilon<1$, there exist $\widetilde{\mathscr{E}}\in\mathcal{F}_{\mathrm{NN}}^{\mathscr{E}}=\mathcal{F}\left(D,n;L_{\mathscr{E}},p_{\mathscr{E}},K_{\mathscr{E}},\kappa_{\mathscr{E}},R_{\mathscr{E}}\right)$, $\widetilde{\mathbf{A}}\in\mathbb{R}^{n\times n}$ with $\|\widetilde{\mathbf{A}}\|_{2}\leq\rho$ and $\widetilde{\mathscr{D}}\in\mathcal{F}_{\mathrm{NN}}^{\mathscr{D}}=\mathcal{F}\left(n,D;L_{\mathscr{D}},p_{\mathscr{D}},K_{\mathscr{D}},\kappa_{\mathscr{D}},R_{\mathscr{D}}\right)$ such that
@@
By [16, Theorem 2.2], for any $0<\varepsilon<1$, let $\varepsilon_{1}=\frac{\varepsilon}{(1+C_{\phi}\rho\sqrt{n})}<1$ , there exists an encoder network $\widetilde{\mathscr{E}}$ such that
@@
Moreover, by Kirszbraun’s theorem [22], the inverse map $\phi=\varphi^{-1}$ admits a Lipschitz extension (still denoted by $\phi$) to $\Lambda,\Lambda]^{n}$ with Lipschitz constant $C_{\phi}$. Applying the approximation result of [38, Lemma 8] to this extended $\phi$, there exists a decoder network $\widetilde{\mathscr{D}}$ such that
@@
$\chi_{1}=2\mathbb{E}_{\mathcal{S}}$ $\left[\frac{1}{N}\sum_{i=1}^{N}\left(\|\widehat{\mathscr{G}}(\mathbf{x}_{i})-\mathbf{F}(\mathbf{x}_{i})\|_{2}^{2}+\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x}_{i})-\mathbf{x}_{i}\|_{2}^{2}+\|\widehat{\mathscr{H}}(\mathbf{x}_{i})-\widehat{\mathscr{E}}(\mathbf{F}(\mathbf{x}_{i})\|_{2}^{2})\right)\right]$
@@
$\chi_{2}=$ $\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(x,x^{+})\sim\mathbb{P}_{X,X^{+}}}\|\widehat{\mathscr{G}}(\mathbf{x})-\mathbf{F}(\mathbf{x})\|_{2}^{2}-2\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{G}}(\mathbf{x}_{i})-\mathbf{F}(\mathbf{x}_{i})\|_{2}^{2}\right]$
@@
$+\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(x,x^{+})\sim\mathbb{P}_{X,X^{+}}}\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x})-\mathbf{x}\|_{2}^{2}-2\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x}_{i})-\mathbf{x}_{i}\|_{2}^{2}\right]$
@@
(26) $+\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(x,x^{+})\sim\mathbb{P}_{X,X^{+}}}\|\widehat{\mathscr{H}}(\mathbf{x})-\widehat{\mathscr{E}}(\mathbf{x})\|_{2}^{2}-2\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{H}}(\mathbf{x}_{i})-\widehat{\mathscr{E}}(\mathbf{F}(\mathbf{x}_{i})\|_{2}^{2}\right]~.$
@@
Let $\widetilde{\mathscr{E}}$, and $\widetilde{\mathscr{D}}$ be the networks in Lemma A.3 with accuracy $\varepsilon_{1}=\frac{\varepsilon}{1+C_{\phi}\rho\sqrt{n}}<1$. For the first term $2\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum\limits_{i=1}^{N}\|\widehat{\mathscr{G}}(\mathbf{x}_{i})-\mathbf{F}(\mathbf{x}_{i})\|_{2}^{2}\right]$ of $\chi_{1}$, we have
@@
$2\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{G}}(\mathbf{x}_{i})-\mathbf{F}(\mathbf{x}_{i})\|_{2}^{2}\right]$ $\leq 2\mathbb{E}_{\mathcal{S}}\inf_{\mathscr{G}\in\mathcal{F}_{NN}^{\mathscr{G}}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\mathscr{G}(\mathbf{x}_{i})-\mathbf{F}(\mathbf{x}_{i})\|_{2}^{2}\right]$
@@
$\leq 2\inf_{\mathscr{G}\in\mathcal{F}_{NN}^{\mathscr{G}}}\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\mathscr{G}(\mathbf{x}_{i})-\mathbf{F}(\mathbf{x}_{i})\|_{2}^{2}\right]$
@@
$\leq 2\mathbb{E}_{(x,x^{+})\sim\mathbb{P}_{X,X^{+}}}\left[\|\widetilde{\mathscr{D}}\circ{\mathbf{A}^{\star}}\circ\widetilde{\mathscr{E}}(\mathbf{x})-\mathbf{F}(\mathbf{x})\|_{2}^{2}\right]$
@@
$\leq 2D\sup_{\mathbf{x}\in\mathcal{M}}\|\widetilde{\mathscr{D}}\circ{\mathbf{A}^{\star}}\circ\widetilde{\mathscr{E}}(\mathbf{x})-\mathbf{F}(\mathbf{x})\|_{\infty}^{2}$
(27) $\leq 2D\varepsilon^{2}.$
@@
(28) $2\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x}_{i})-\mathbf{x}_{i}\|_{2}^{2}\right]\leq 2D\varepsilon^{2}.$
@@
$2\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{H}}(\mathbf{x}_{i})-\widehat{\mathscr{E}}(\mathbf{F}(\mathbf{x}_{i}))\|_{2}^{2}\right]$
$\leq$ $2\inf_{\mathscr{H}\in\mathcal{F}_{AE}^{\mathscr{H}}}\mathbb{E}_{(x,x^{+})\sim\mathbb{P}_{X,X^{+}}}\left[\|\mathscr{H}(\mathbf{x})-{\mathscr{E}}(\mathbf{F}(\mathbf{x}))\|_{2}^{2}\right].$
We next bound the right-hand side by evaluating it at the candidate $\widetilde{\mathscr{H}}:=\widetilde{\mathbf{A}}\,\widetilde{\mathscr{E}}$. For any $\mathbf{x}\in\mathcal{M}$, by the triangle inequality,

### 29 Apr 2026 v2

@@ 5.3 Lorenz–96 model
where $\mathcal{H}$ denotes the observation operator. We consider both linear and nonlinear observation settings. In the linear observation setting, the observation operator $\mathcal{H}$ is chosen as a subsampling operator that extracts every other state variable,
@@
with observation noise level $\sigma=1$. This setup represents a partially observed scenario in which only half of the state variables are directly measured.
Dense observation regime
We first consider a relatively dense observation setting with observations available every $\Delta t=0.1$. LAE models are trained with latent dimensions $n\in\{64,80,100\}$, and the resulting LAE-EnKF is compared with AE-EnKF, DAE-EnKF, and an EnKF baseline. Tab. 3 reports the global relative RMSE $\mathrm{E}_{\mathrm{Rel},1:K}$ over the full state for all methods, both with and without localization. Localization significantly improves the performance of EnKF and AE-EnKF, consistent with the well-known limitation of EnKF for the Lorenz–96 system. Without localization, sampling errors in the empirical covariance introduce spurious long-range correlations, leading to instability and large reconstruction errors. Since EnKF and AE-EnKF operate in the original state space, localization is essential for stabilizing their performance. Accordingly, unless otherwise specified, all references to EnKF and AE-EnKF hereafter correspond to their localized implementations. In contrast, DAE-EnKF and LAE-EnKF operate in a latent space, where spurious correlations are inherently reduced. Localization provides negligible improvement for these methods. In particular, LAE-EnKF achieves low errors, with nearly identical performance with and without localization, indicating that localization is unnecessary in this setting. This indicates that latent representations mitigate sampling errors through structured dynamics.
Table 3: Relative RMSE $\mathrm{E}_{\mathrm{Rel},1:K}$ for different methods with varying latent dimension $n$, reported without (with) covariance localization, under two observation intervals.
$\Delta t$ $\boldsymbol{n}$ EnKF AE-EnKF DAE-EnKF LAE-EnKF
0.5977 (0.1590)
$64$ 0.4504 (0.2027) 0.8276 (0.8195) 0.1475 (0.1480)
$0.1$ $80$ 0.3793 (0.2026) 0.8202 (0.8697) 0.1509 (0.1512)
$100$ 0.3563 (0.2080) 0.6696 (0.6804) 0.1426 (0.1431)
0.3223 (0.1689)
$64$ 0.4376 (0.2416) 1.1994 (1.0393) 0.1862 (0.1864)
$0.2$ $80$ 0.3502 (0.2334) 0.7244 (0.7274) 0.1704 (0.1708)
$100$ 0.2723 (0.2360) 0.6567 (0.6452) 0.1679 (0.1682)
Fig. 13 shows the reconstructed spatial fields and the corresponding errors for the representative case $n=64$, while Fig. 14 presents the temporal evolution of one observed and one unobserved state variable. DAE-EnKF exhibits larger reconstruction errors and irregular temporal behavior, indicating instability in the learned latent dynamics. In contrast, LAE-EnKF consistently achieves lower errors and smoother temporal evolution for both observed and unobserved variables. This demonstrates that enforcing stable linear dynamics in the latent space substantially improves the recovery of hidden states in a chaotic system.
@@
Figure 13: Reconstructed spatial state fields (Left) and corresponding error (Right) with the observation interval $\Delta t=0.1$ and latent dimension $n=64$. From top to bottom, the results show the localization EnKF, localization AE-EnKF, DAE-EnKF and the LAE-EnKF.
@@
Figure 14: Time evolution of a representative observed variable (Left) and an unobserved variable (Right) reconstructed with observation interval $\Delta t=0.1$ and latent dimension $n=64$, compared with the corresponding true trajectories. From top to bottom, the results show the localization EnKF, localization AE-EnKF, DAE-EnKF and the LAE-EnKF.
Sparse observation regime
@@
Figure 15: Reconstructed spatial state fields (Left) and corresponding error (Right) with the observation interval $\Delta t=0.2$ and latent dimension $n=64$. From top to bottom, the results show the localization EnKF, localization AE-EnKF, DAE-EnKF and the LAE-EnKF.
@@
Nonlinear observation setting
We further consider a nonlinear observation operator defined as
$\mathcal{H}(\mathbf{x})=\arctan(\mathbf{x}).$
Under nonlinear observation settings, we compare the localized EnKF with the proposed LAE-EnKF. For LAE-EnKF, we investigate the impact of different latent observation operators $\mathbf{H}$, including the identity operator, random operators, and learned operators. Fig. 14 shows the reconstructed trajectory of a representative state component, with the corresponding global RMSE $\mathrm{E}_{\mathrm{Rel},1:K}$ annotated above each figure. The results show that the performance of the standard EnKF deteriorates under nonlinear observations, whereas LAE-EnKF remains stable and achieves more accurate state reconstruction, closely matching the true trajectory. Moreover, structured choices of $\mathbf{H}$, such as the identity or random operators, generally outperform learned operators in terms of reconstruction accuracy. This behavior may be attributed to the fact that structured operators are more likely to preserve the geometric and covariance structure of the latent space, while a fully learned $\mathbf{H}$ may introduce distortions in dominant dynamical directions, thereby affecting the effectiveness of the Kalman update.
[…]
Figure 17: Reconstruction of representative state components under nonlinear observations. Comparison between localized EnKF and LAE-EnKF under different latent observation operators, including the identity, random, and learned operators. The global RMSE $\mathrm{E}_{\text{Rel},1:K}$ is reported above each panel.
@@ 6 Conclusion
We denote by $\mathcal{F}_{\mathrm{NN}}(d_{1},d_{2};L,p,K,\kappa,R)$ the class of neural networks introduced in [31], where $d_{1}$ and $d_{2}$ are the input and output dimensions. The networks have depth $L$, width $p$, and at most $K$ nonzero parameters, each bounded by $\kappa$ in magnitude. The effective parameter count $K$ may depend on $(d_{1},d_{2})$. For inputs satisfying $\|\mathbf{x}\|_{\infty}\leq B$, the network outputs are uniformly bounded by $R$. The proof is based on controlling both approximation and estimation errors through a covering number bound for the class $\mathcal{F}_{\mathrm{NN}}$ under the uniform norm. To make the argument precise, we first introduce the assumptions and auxiliary lemmas needed for the subsequent analysis. For the subsequent analysis, we use the mixed $L^{\infty,\infty}$ norm defined by $\|\mathbf{f}\|_{L^{\infty,\infty}}:=\sup_{\mathbf{x}\in\Omega}\|\mathbf{f}(\mathbf{x})\|_{\infty}$. The latent variables are assumed to lie in a bounded domain, taken as $\Lambda,\Lambda]^{n}$ up to approximation of the manifold parametrization $\varphi$.
Definition A.1 (Covering number [45, Definition 2.1.5]).
@@
Suppose Assumption 1 holds. For any $0<\varepsilon<1$, there exist $\widetilde{\mathscr{E}}\in\mathcal{F}_{\mathrm{NN}}^{\mathscr{E}}=\mathcal{F}\left(D,n;L_{\mathscr{E}},p_{\mathscr{E}},K_{\mathscr{E}},\kappa_{\mathscr{E}},R_{\mathscr{E}}\right)$, $\widetilde{\mathbf{A}}\in\mathbb{R}^{n\times n}$ with $\|\widetilde{\mathbf{A}}\|_{2}\leq\rho{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ (\rho\leq 1)}$ and $\widetilde{\mathscr{D}}\in\mathcal{F}_{\mathrm{NN}}^{\mathscr{D}}=\mathcal{F}\left(n,D;L_{\mathscr{D}},p_{\mathscr{D}},K_{\mathscr{D}},\kappa_{\mathscr{D}},R_{\mathscr{D}}\right)$ such that
@@
By [20, Theorem 2.2], for any $0<\varepsilon<1$, let $\varepsilon_{1}=\frac{\varepsilon}{(1+C_{\phi}\rho\sqrt{n})}<1$ , there exists an encoder network $\widetilde{\mathscr{E}}$ such that
@@
Moreover, by Kirszbraun’s theorem [27], the inverse map $\phi=\varphi^{-1}$ admits a Lipschitz extension (still denoted by $\phi$) to $\Lambda,\Lambda]^{n}$ with Lipschitz constant $C_{\phi}$. Applying the approximation result of [48, Lemma 8] to this extended $\phi$, there exists a decoder network $\widetilde{\mathscr{D}}$ such that
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(\mathbf{x},\mathbf{x}^{+})\sim\mathbb{P}_{X,X^{+}}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Bigl[\|\widehat{\mathscr{G}}(\mathbf{x})-\mathbf{x}^{+}\|_{2}^{2}+\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x})-\mathbf{x}\|_{2}^{2}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\|\widehat{\mathscr{H}}(\mathbf{x})-\widehat{\mathscr{E}}(\mathbf{x}^{+})\|_{2}^{2}+R(\widehat{\mathbf{A}})\Bigr]}$
@@
[…]
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\,\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(\mathbf{x},\mathbf{x}^{+})\sim\mathbb{P}_{X,X^{+}}}\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x})-\mathbf{x}\|_{2}^{2}-C_{1}\,\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x}_{i})-\mathbf{x}_{i}\|_{2}^{2}\right]}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\,\mathbb{E}_{\mathcal{S}}\mathbb{E}_{(\mathbf{x},\mathbf{x}^{+})\sim\mathbb{P}_{X,X^{+}}}\|\widehat{\mathscr{H}}(\mathbf{x})-\widehat{\mathscr{E}}(\mathbf{x}^{+})\|_{2}^{2}-\,C_{1}\,\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{H}}(\mathbf{x}_{i})-\widehat{\mathscr{E}}(\mathbf{x}_{i}^{+})\|_{2}^{2}\right].}$
@@
Let $\widetilde{\mathscr{E}}$, and $\widetilde{\mathscr{D}}$ be the networks in Lemma A.3 with accuracy $\varepsilon_{1}=\frac{\varepsilon}{1+C_{\phi}\rho\sqrt{n}}<1$. For the first term ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum\limits_{i=1}^{N}\|\widehat{\mathscr{G}}(\mathbf{x}_{i})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}_{i}^{+}}\|_{2}^{2}\right]$ of $\chi_{1}$, we have
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\,\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{G}}(\mathbf{x}_{i})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}_{i}^{+}}\|_{2}^{2}\right]$
@@
$\leq$ $\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\,\mathbb{E}_{\mathcal{S}}\inf_{\mathscr{G}\in\mathcal{F}_{NN}^{\mathscr{G}}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\mathscr{G}(\mathbf{x}_{i})-\mathbf{F}(\mathbf{x}_{i})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{w}_{i}}\|_{2}^{2}\right]$
$\leq$ $\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\inf_{\mathscr{G}\in\mathcal{F}_{NN}^{\mathscr{G}}}\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\mathscr{G}(\mathbf{x}_{i})-\mathbf{F}(\mathbf{x}_{i})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{w}_{i}}\|_{2}^{2}\right]$
@@
$\leq$ $\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\,\mathbb{E}_{(\mathbf{x},\mathbf{x}^{+})\sim\mathbb{P}_{X,X^{+}}}\left[\|\widetilde{\mathscr{D}}\circ{\mathbf{A}^{\star}}\circ\widetilde{\mathscr{E}}(\mathbf{x})-\mathbf{F}(\mathbf{x})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{w}}\|_{2}^{2}\right]$
@@
$\leq$ $\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\,D\sup_{\mathbf{x}\in\mathcal{M}}\|\widetilde{\mathscr{D}}\circ{\mathbf{A}^{\star}}\circ\widetilde{\mathscr{E}}(\mathbf{x})-\mathbf{F}(\mathbf{x})\|_{\infty}^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}\mathbb{E}\|\mathbf{w}\|_{2}^{2}}$
(23) $\leq$ $\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\,D\,\varepsilon^{2}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}\,\mathbb{E}\|\mathbf{w}\|_{2}^{2}}.$
@@
(24) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\,\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{D}}\circ\widehat{\mathscr{E}}(\mathbf{x}_{i})-\mathbf{x}_{i}\|_{2}^{2}\right]\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\,D\,\varepsilon^{2}.$
@@
$\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\,\mathbb{E}_{\mathcal{S}}\left[\frac{1}{N}\sum_{i=1}^{N}\|\widehat{\mathscr{H}}(\mathbf{x}_{i})-\widehat{\mathscr{E}}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{x}_{i}^{+}})\|_{2}^{2}\right]$
$\leq$ $\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}}\inf_{\mathscr{H}\in\mathcal{F}_{AE}^{\mathscr{H}}}\mathbb{E}_{(\mathbf{x},\mathbf{x}^{+})\sim\mathbb{P}_{X,X^{+}}}\left[\|\mathscr{H}(\mathbf{x})-{\mathscr{E}}\left(\mathbf{F}(\mathbf{x})+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{w}}\right)\|_{2}^{2}\right]$
$\leq$ $\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{1}\inf_{\mathscr{H}\in\mathcal{F}_{AE}^{\mathscr{H}}}\mathbb{E}_{(\mathbf{x},\mathbf{x}^{+})\sim\mathbb{P}_{X,X^{+}}}\bigg\{\frac{C_{1}}{C_{1}-1}\left[\|\mathscr{H}(\mathbf{x})-{\mathscr{E}}(\mathbf{F}(\mathbf{x}))\|_{2}^{2}\right]}$
$\qquad\qquad\qquad\qquad\qquad\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\,C_{1}\|{\mathscr{E}}(\mathbf{F}(\mathbf{x}))-\mathscr{E}(\mathbf{F}(\mathbf{x})+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbf{w}})\|_{2}^{2}\bigg\}.}$
We next bound the first row of last step by evaluating it at the candidate $\widetilde{\mathscr{H}}:=\widetilde{\mathbf{A}}\,\widetilde{\mathscr{E}}$. For any $\mathbf{x}\in\mathcal{M}$, by the triangle inequality,
@@
For the second row, for any $\mathbf{x}\in\mathcal{M}$, by the triangle inequality,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\widetilde{{\mathscr{E}}}(\mathbf{F}(\mathbf{x}))-\widetilde{\mathscr{E}}(\mathbf{F}(\mathbf{x})+\mathbf{w})\|_{2}}$
[…]
$\quad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}+\,\|\varphi(\mathbf{F}(\mathbf{x})+\mathbf{w})-\widetilde{\mathscr{E}}(\mathbf{F}(\mathbf{x})+\mathbf{w})\|_{2}}$
[…]